Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Sarah
Log entry time 19:37:36 on September 18, 2009
Entry number 290402
keyword=Run Coordinator's Report for September 17 Swing to September 18 Day
Run Coordinator's Report for September 17 Swing to September 18 Day
Summary
We took production data and Compton data throughout this period. A
position scan was also done during the owl shift to investigate whether
we are hitting the aluminum and causing target boiling.
Significant News
- Last night's swing shift started with us switching back to
production running from the Moller measurement that ended at the end of
the previous shift. The beam was tuned, the spot size verified as okay
with harp scans, the raster was verified, and production running began.
The shift took production data throughout the night, with the only snag
being a reboot of the target ioc. After reaching 1M good pairs, the
half-wave plate was changed, but flipped did not work, so the values had
to be put in by hand (halog
entry 290216).
- The scaler is fixed (maybe!). See halog
entry 290177.
- A beam position scan was done during the owl shift to see if we are
hitting the aluminum and causing target boiling that way. See halog
entry 290295 and halog
entry 290297.
- After the scan, we went back to production running throughout the
rest of the owl and day shifts. There was some time lost to CODA reboots
and IOC reboots (by MCC and our target IOC).
Plans
- The detailed run plan for the night is in halog
290400.
- Take lots of production data! Watch the lumi widths to see if
anything changes. Please check the histograms on the online data monitor
with the reference ones on the shift how-to page (http://hallaweb.jlab.org/experiment/HAPPEX/opmanual/opmanual.html#HPXStartingonlineanalysis).
- Run spot++ every couple of hours and log it. Change the prescale
factor PS8 to 1 and PS1 and 2 to 99999.
- Take Compton data, starting a new run every two hours. Call the
Compton experts if the background rates get so high that they
consistently trip off the Compton detector HV.
- After an extended period when the beam is down, establish good beam
(do harps, spot++, check raster), then go to production running.
Concerns
- Flipper needs to be fixed!
- Scot Spiegel would like to make an access if we are going to be
down for a while at a time when we can go to controlled access (in the
daytime, 8am-5pm) to get some poly for shielding. If such a time
presents itself, his extension is x5900, his cell is 757-536-5900, and
his pager is 757-584-5900.
- Page Heidi if there is an access in the daytime (6915).
- If anyone must reset a magnet, they must make a halog entry
stating which magnet on which arm, and what was wrong with it to require
a reset. The word "tech" must appear in the subject line.
- Be sure that the prompt analysis is done for every run!
- Please log CODA reboots in the halog and shift summary so that it
is easier to track how much time is lost to these reboots.
- The BTA program doesn't subtract any time to speak of for
beam trips. Please manually subtract a full minute for each beam trip in
the beam time accounting.
- Please write a detailed shift summary (including runs taken and any
activities) and submit the BTA for each shift.
- Please make sure that the shift checklist is done once each shift.
- If you have to page or call someone for support, please do so from a
counting house phone or it will be ignored!
- Keep a close eye on the lumi widths as we take data.
- Please continue to take runs even if there is no beam. These data
can be used to determine pedestals. Please label these runs as having no
beam.
- The PVDIS people would like their scintillators on. See halog
entry 290362.
- There are lots of empty shifts that need to be filled immediately!
Please sign up!
- Is bpm12 still saturating?
- Hall C charge asymmetry feedback.
- The dithering looks non-invertible. Someone should look at this.
- We need to do position scans to make sure that we are not hitting
the aluminum.
- The electronics pickup data needs to be looked at carefully
(pedestals and Pockels cell off data).
- MCC has advised us to tell them when we want a Moller measurement
before we take our target out so that they can do their all save and
everything with things as they are.
Coulombs So Far
- Coulombs taken in this time period: 6.7 C
- Coulombs taken in all so far: 50.7 C as of 14:14pm
Shift Summaries and BTA Reports